A 500.0 mL vessel contain 1.35g chlorine gas has a pressure of 760 mmHg, Oxygen gas added gas pressure rises to 865 mmHg. Assume gases don't react each other and temperature remains constant, calculate mass of oxygen gas added into the vessel

1
Expert's answer
2021-11-02T02:23:45-0400

P1/M1=P2/M2

760/1.35=865/M2

562.96= 865/M2

M2= 865/562.96

= 1.54g